OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 4x - y - 3z = -3 x - 2y - 2z = -3 x + y + 3z = -2

OpenStudy (aravindg):

solve 2 eqns at once

OpenStudy (anonymous):

I don't understand how to do that..

OpenStudy (aravindg):

i mean solve the first two eqns to eliminate one variable

OpenStudy (anonymous):

set up a 3x3 matrix and use Gaussian Elimination with backward sustitution

OpenStudy (anonymous):

x=-13/4 y=16/7 z=-1